OpenStudy (anonymous):

Measurement unit Relationship flop 4 flops = 3 hats bit 2 bits = 0.4 den hat den 1.7 dens = 5 hats Answer: 7 bits = ____ hats 3 flops = ___bits

OpenStudy (austinl):

This will be similar to dimensional analysis. Do you have any thoughts?

OpenStudy (anonymous):

Not quire sure how to set it up

OpenStudy (ranga):

7 bits = ____ hats We need to connect bits and hats Let us look for which two equations will help connect dens and hats: 2 bits = 0.4 den ------ (1) 1.7 dens = 5 hats ------ (2) From (2) you can find 1 den equals how many hats Put that in (1) and you will have 2 bits equals how many dens Then find 7 bits equals how many dens

OpenStudy (anonymous):

1.4

OpenStudy (anonymous):

Well, no thats wrong.

OpenStudy (ranga):

Can you show the work that got you 1.4?

OpenStudy (anonymous):

7/5

OpenStudy (ranga):

2 bits = 0.4 den ------ (1) 1.7 dens = 5 hats ------ (2) Divide both sides of (2) by 1.7: 1 den = 5/1.7 hats 1 den = 2.94 hats. Multiply both sides by 0.4 1 * 0.4den = 2.94 * 0.4 hats 0.4 den = 1.176 hats Put this in (1) 2 bits = 1.176 hats divide both sides by 2 1 bit = .588 hat multiply both sides by 7 7 bits = .588 * 7 = 4.116 hats If you round it to two decimals you get 4.12 or round it to one decimal you get 4.1 hat 7 bits = 4.1 hat Follow similar procedure for the other question.

OpenStudy (anonymous):

So what equation would be used to find this answer besides 4 flops= 3hats? Would it be 7 bits=4.1 hats?

OpenStudy (ranga):

All three of them: 4 flops = 3 hats connects F & H 2 bits = 0.4 den connects B & D 1.7 dens = 5 hats connects D & H For the second question you want to connect F & B

OpenStudy (anonymous):

I am completely lost.